# Debate, Train, Evolve: Self-Evolution of Language Model Reasoning

### v1

@@ 4.2 Main Results
Our main results are organized into three main parts: (1) First, we evaluate the effectiveness of Debate–Train–Evolve (DTE) framework, (2) Next, we test its generalization across different reasoning tasks by transferring evolved models to new datasets, and (3) Finally, we analyze the extent of model self-evolution through iterative rounds.
@@
Evolved model using DTE shows an average gain of 8.92% accuracy on GSM-PLUS compared to its vanilla performance. Table 1 contrasts three settings: the single base model (“Original”), a three-agent debate with our RCR prompt (“MAD”), and the evolved single model produced by one Debate–Train–Evolve pass. On GSM-Plus—the hard math dataset—DTE improves every model, with an average gain of +2.38 points over three-agent MAD. Qwen-1.5B shows the largest jump (+13.92 pts), confirming that evolution is most helpful when the base model has head-room and the debate provides diverse traces. On GSM8K the average gain is smaller ( +0.84 pts) because several models were already near their ceiling after debate. ARC-Challenge sees a mixed results: large models benefit (+3.67 pts for Qwen-7B, +8.88 pts for Llama-8B) while small models drift by pt. Overall, DTE shows a mean improvement of 3.06 pts over single model and +1.09 pts over MAD while restoring single-pass inference.
@@ 4.3 Ablation Studies
We observe two consistent trends here: First, when the individual agents have comparable standalone accuracy, cross-family mixtures beat homogeneous agents team, supporting the idea that architectural diversity yields complementary reasoning paths. Second, when the pool mixes a strong and a weaker model, the debate result gravitates toward the stronger member—adding the weaker agent neither helps nor seriously harms, suggesting that diversity only helps when all agents can contribute novel insights. Complete results for every dataset and roster is available in Appendix B.
@@
Figure 5 plots GSM-Plus accuracy as we grow the number of GRPO training steps from 2K to 10K. All models share the similiar trend: rapid gains up to about 8K steps followed by saturation. Small and mid-size models profit the most from the early updates—Qwen-1.5B climbs 8.0 pts between 2K and 6K samples—whereas larger models such as Qwen-14B rise more slowly but steady. Beyond 8K the curve flattens: the average improvement from 8K 10 k is only +0.32 pts while wall-clock time grows by 25%.
@@
Figure 6 plots GSM8K and GSM-Plus accuracy for Qwen-1.5B after the first and second evolution rounds under four sampling temperatures. When we keep the original exploratory setting () the model loses 2.0 pts on GSM8K and gains only 13.5 pts on GSM-Plus—well below the +33.5 pts it achieved in Round 1—confirming a clear case of catastrophic forgetting. Lowering the temperature stabilises training: at Round-2 accuracy is within 0.9 pts of Round 1 on GSM-Plus and almost fully recovers on GSM8K; a deterministic schedule () even adds +3.3 pts on GSM8K but plateaus on GSM-Plus.
@@ 5 Conclusion
This study explore the self-evolution of language models using publicly available benchmarks and datasets such as GSM8K, ARC, and CommonsenseQA. All data used in our experiments are non-sensitive and freely accessible, ensuring compliance with ethical research standards and reproducibility. Our method involves fine-tuning on model-generated content, without introducing or relying on any human-annotated private data.
Acknowledgements
This work was supported by NSF NAIRR Pilot with PSC Neocortex, NCSA Delta; Amazon, Cisco Research, Commonwealth Cyber Initiative, Amazon–Virginia Tech Center for Efficient and Robust Machine Learning, and Sanghani Center for AI and Data Analytics at Virginia Tech. The views, findings, conclusions, and recommendations expressed in this work are those of the authors and do not necessarily reflect the opinions of the funding agencies.
@@ Appendix A Datasets Details
We evaluate our Multi-Agent Debate (MAD) approach on five diverse reasoning benchmarks. In the following, we briefly describe each dataset along with their splits. In this paper, we use the test split to evaluate all Small Language Models (SLMs). Table 5 summarizes the splits for each dataset.
@@
GSM8K
Cobbe et al. (2021) is a collection of high-quality grade school math word problems that require multi-step reasoning. In the main configuration, the dataset contains a total of 8,790 examples, with 7,473 examples in the training split and 1,319 examples in the test split. These problems typically require between 2 and 8 steps to solve, making it an excellent benchmark for evaluating mathematical reasoning capabilities.
GSM-Plus
Li et al. (2024) extends the GSM8K benchmark with more challenging and diverse mathematical word problems. GSM+ problems generally require more sophisticated multi-step reasoning and often involve more complex mathematical concepts than those in the original GSM8K dataset.
ARC
Clark et al. (2018) comprises two subsets of multiple-choice science questions:
ARC-Easy: Contains 2,251 train, 570 validation, and 2,376 test examples. These questions are answerable by most middle school students and test basic science knowledge.
ARC-Challenge: Contains 1,119 train, 299 validation, and 1,172 test examples. These questions are more challenging and typically answered incorrectly by both retrieval-based algorithms and word co-occurrence algorithms.
CommonsenseQA
Talmor et al. (2019) requires commonsense reasoning to answer multiple-choice questions. It has 9,741 training examples, 1,221 validation examples, and 1,140 test examples. The questions are specifically designed to test commonsense knowledge and reasoning capabilities that go beyond simple factual recall.
@@ Appendix B Implementation Details
Training
For model fine-tuning, we used GRPO to enhance Language Models on the targeted reasoning tasks. Our training pipeline utilized the Unsloth 111https://github.com/unslothai/unsloth and TRL 222https://github.com/huggingface/trl libraries for efficient parameter-efficient fine-tuning with QLoRA Dettmers et al. (2023). Models were trained with a LoRA Hu et al. (2021) rank of 128 and target modules including query, key, value, output, gate, up, and down projections. We used 8-bit Adam optimizer with beta parameters of (0.9, 0.99) and a weight decay of 0.1. The learning rate was set to 5e-6 with a cosine decay schedule and 10% warmup ratio. Training proceeded for 10,000 steps with a per-device batch size of 8. To improve output formatting, we implemented a multi-component reward function consisting of: (1) an answer correctness reward, (2) format adherence rewards for XML tags structure, (3) a numeric response reward, and (4) a tag-counting reward to incentivize proper tag usage. Each model was instructed to output responses in a structured XML format with separate <reasoning> and <answer> tags to facilitate consistent answer extraction and evaluation. To manage memory constraints on high-end GPUs, we set the maximum sequence length to 2048 tokens with a 512-token maximum for prompts and 1536 tokens for model completions.
Inference
We conducted all model inferences using NVIDIA H100-80GB, A100-80GB, L40-48GB, and A40-48GB GPUs. For efficient inference, we used the vLLM library Kwon et al. (2023) 333https://docs.vllm.ai/en/latest/, dynamically allocating the required number of GPUs to load each model. Multi-GPU utilization was enabled using Hugging Face Accelerate 444https://github.com/huggingface/accelerate for model sharding and speed optimization.

### v2

@@ 4.2 Main Results
Our main results are organized into three main parts: (1) First, we evaluate the effectiveness of DTE framework, (2) Next, we test its generalization across different reasoning tasks, and (3) Finally, we analyze the extent of model self-evolution through iterative rounds.
@@
Evolved model using DTE shows an average gain of 8.92% accuracy on GSM-PLUS compared to its vanilla performance. Table 1 contrasts three settings: the single base model (“Original”), a three-agent debate with our RCR prompt (“MAD”), and the evolved single model produced by one Debate–Train–Evolve pass. On GSM-Plus-the hard math dataset-DTE improves every model, with an average gain of +2.38 points over three-agent MAD. Qwen-1.5B shows the largest jump (+13.92 pts), confirming that evolution is most helpful when the base model has head-room and the debate provides diverse traces. On GSM8K the average gain is smaller ( +0.84 pts) because several models were already near their ceiling after debate. ARC-Challenge sees a mixed results: large models benefit (+3.67 pts for Qwen-7B, +8.88 pts for Llama-8B) while small models drift by pt. Overall, DTE shows a mean improvement of 3.06 pts over single model and +1.09 pts over MAD while restoring single-pass inference.
@@ 4.3 Ablation Studies
We observe two consistent trends here: First, when the individual agents have comparable standalone accuracy, cross-family mixtures beat homogeneous agents team, supporting the idea that architectural diversity yields complementary reasoning paths. Second, when the pool mixes a strong and a weaker model, the debate result gravitates toward the stronger member-adding the weaker agent neither helps nor seriously harms, suggesting that diversity only helps when all agents can contribute novel insights. Complete results for every dataset and roster is available in Appendix B.
@@
Figure 5 plots GSM-Plus accuracy as we grow the number of GRPO training steps from 2K to 10K. All models share the similiar trend: rapid gains up to about 8K steps followed by saturation. Small and mid-size models profit the most from the early updates-Qwen-1.5B climbs 8.0 pts between 2K and 6K samples-whereas larger models such as Qwen-14B rise more slowly but steady. Beyond 8K the curve flattens: the average improvement from 8K 10 k is only +0.32 pts while wall-clock time grows by 25%.
@@
Figure 6 plots GSM8K and GSM-Plus accuracy for Qwen-1.5B after the first and second evolution rounds under four sampling temperatures. When we keep the original exploratory setting () the model loses 2.0 pts on GSM8K and gains only 13.5 pts on GSM-Plus-well below the +33.5 pts it achieved in Round 1-confirming a clear case of catastrophic forgetting. Lowering the temperature stabilises training: at Round-2 accuracy is within 0.9 pts of Round 1 on GSM-Plus and almost fully recovers on GSM8K; a deterministic schedule () even adds +3.3 pts on GSM8K but plateaus on GSM-Plus.
@@ 5 Conclusion
Acknowledgements
This work was supported by the NSF #2442253, NSF NAIRR Pilot with PSC Neocortex and NCSA Delta, Cisco Research, NVIDIA, Amazon, the Commonwealth Cyber Initiative, the Amazon–Virginia Tech Center for Efficient and Robust Machine Learning, the Sanghani Center for AI and Data Analytics at Virginia Tech, the Virginia Tech Innovation Campus, Children’s National Hospital, and the Fralin Biomedical Research Institute at Virginia Tech. The views, findings, conclusions, and recommendations expressed in this work are those of the authors and do not necessarily reflect the opinions of the funding agencies.
@@ Acknowledgements
This study explore the self-evolution of language models using publicly available benchmarks and datasets such as GSM8K, MATH, ARC, GPQA, and CommonsenseQA. All data used in our experiments are non-sensitive and freely accessible, ensuring compliance with ethical research standards and reproducibility. Our method involves fine-tuning on model-generated content, without introducing or relying on any human-annotated private data.
AI Assistance:
We used ChatGPT assistance for parts of the Appendix, such as generating LaTeX code for tables and refining text written by the authors. All AI-generated content was carefully reviewed and revised by the authors to ensure accuracy and clarity.
@@ Appendix A Datasets Details
We evaluate our approach on seven diverse reasoning benchmarks that test different aspects of model capabilities. Each dataset was chosen to provide complementary challenges in reasoning tasks. Table 5 summarizes the dataset statistics.
@@
GSM8K (Cobbe et al., 2021) contains 8,790 grade school math word problems requiring multi-step reasoning. We use 7,473 training examples and evaluate on 1,319 test problems. Each problem needs 2-8 reasoning steps to solve.
GSM-Plus (Li et al., 2024) provides 2,400 adversarial variations of GSM8K problems designed to test robustness. These problems include more complex numerical values and additional reasoning steps compared to the original dataset.
MATH (Hendrycks et al., 2021) consists of 12,500 competition mathematics problems from AMC 10, AMC 12, AIME, and other competitions. Problems span topics from algebra to calculus with difficulty levels from 1 to 5. We use the standard splits of 7,500 training and 5,000 test problems.
ARC (Clark et al., 2018) includes science questions at two difficulty levels. ARC-Easy has 2,251 training, 570 validation, and 2,376 test questions answerable by middle school students. ARC-Challenge contains 1,119 training, 299 validation, and 1,172 test questions that are challenging for retrieval-based methods.
GPQA Main (Rein et al., 2024) presents 448 graduate-level multiple-choice questions in biology, physics, and chemistry. These expert-written questions are designed to be “Google-proof”- skilled non-experts achieve only 34% accuracy despite unrestricted web access. We use this as a test-only benchmark.
CommonsenseQA (Talmor et al., 2019) requires commonsense reasoning with 9,741 training, 1,221 validation, and 1,140 test questions. Questions test knowledge that goes beyond factual recall.
@@ Appendix B Implementation Details
Training Setup. We implement GRPO training using the Unsloth222https://github.com/unslothai/unsloth and TRL333https://github.com/huggingface/trl libraries for efficient parameter-efficient fine-tuning. We apply QLoRA (Dettmers et al., 2023) with rank 128 to attention and feed-forward modules (query, key, value, output, gate, up, down projections). Training uses 8-bit AdamW optimization with =0.9, =0.99, weight decay 0.1, and learning rate with cosine decay and 10% warmup. We train for 10,000 steps with batch size 8.
Reward Function. We design a multi-component reward to encourage both correct answers and proper formatting: (1) answer correctness reward with weight 2.0, (2) XML format adherence reward with weight 0.5, (3) numeric response reward with weight 0.5, and (4) tag-counting reward with weight 0.5. Models output structured responses using <reasoning> and <answer> XML tags for consistent evaluation.
Computational Resources. Training runs on NVIDIA H100 (80GB), A100 (80GB), L40 (48GB), and A40 (48GB) GPUs. A single evolution round for a 7B model takes approximately 68 hours on one A100 GPU, consuming about 9600 GPU-hours total. Larger models scale near-linearly with parameter count.
Inference Setup. We use vLLM (Kwon et al., 2023)444https://docs.vllm.ai/en/latest/ for efficient inference with dynamic GPU allocation. Multi-GPU setups use Hugging Face Accelerate555https://github.com/huggingface/accelerate for model sharding and optimization. During debate, we sample at temperature 1.0 for exploration or 0.0 for deterministic responses.
Software and Licenses. All experiments use open-source software. Unsloth and TRL are released under Apache 2.0 license. vLLM uses Apache 2.0 license. All datasets are publicly available with appropriate licenses for research use: GSM8K (MIT), ARC (CC BY-SA 4.0), CommonsenseQA (MIT), MATH (MIT), GSM-Plus (Apache 2.0), and GPQA (available for research with usage restrictions to prevent leakage). Our code and model checkpoints will be released under Apache 2.0 license.
Hyperparameter Selection. We selected hyperparameters through preliminary experiments on validation sets. Key GRPO parameters include clipping threshold =0.2, KL coefficient =0.02, and group size =8. These values balance exploration with training stability across model sizes.